[Hallb] Run status update

Stepan Stepanyan stepanya at jlab.org
Tue Oct 23 08:58:42 EDT 2018

Dear all,

We came back from the beam studies yesterday around 4:30pm, quite smooth 
beam restoration. Took a Moller run with 1/2 wave plate IN, polarization 
54.8+/-1.5%, plot on the run wiki has been updated. Then from 7pm we had 
very nice beam and took production data till 8:30am today. Planned RF 
recovery will be very short, only 2 hours, so we are not planning to 
make an access to the hall unless it is critically necessary. After beam 
is restored, expect some time around 10:30 am, we will continue 
production running. Note, last night we reverted to 2 ph.e. threshold 
trigger setting for HTCC since the 1 ph.e. cut had the same dips in the 
electron phi-distributions (see log entries).

For tomorrow and Thursday situation is not clear. Looks like accelerator 
is eager to provide beam to everyone in parallel to Hall-A's beam work. 
Will know more about details of that two days some time at 2:30 pm. 
There are number of things we want to do that require magnets to be OFF 
and also short restricted access to remove stuff from hall. I will send 
another email when machine plans for Wednesday and Thursday will be more 

As of now, all shifts are ON, we are planning to man the counting house 
for Wednesday and Thursday.

Regards, Stepan

More information about the Hallb mailing list